# HL-IK: A Lightweight Implementation of Human-Like Inverse Kinematics in Humanoid Arms

### v2

@@ Abstract
Traditional IK methods for redundant humanoid manipulators emphasize end-effector (EE) tracking, frequently producing configurations that are valid mechanically but not human-like. We present Human-Like Inverse Kinematics (HL-IK), a lightweight IK framework that preserves EE tracking while shaping whole-arm configurations to appear human-like—without full-body sensing at runtime. The key idea is a learned elbow prior: using large-scale human motion data retargeted to the robot, we train a FiLM-modulated spatio-temporal attention network (FiSTA) to predict the next-step elbow pose from the EE target and a short history of EE–elbow states. This prediction is incorporated as a small residual alongside EE and smoothness terms in a standard Levenberg–Marquardt optimizer, making HL-IK a drop-in addition to numerical IK stacks. Over 183k simulation steps, HL-IK reduces arm-similarity position and direction error by 30.6% and 35.4% on average, and by 42.2% and 47.4% on the most challenging trajectories. Hardware teleoperation on a robot distinct from simulation further confirms the gains in anthropomorphism. HL-IK is simple to integrate, adaptable across platforms via our pipeline, and adds minimal computation, enabling human-like motions for humanoid robots. Project page: https://hl-ik.github.io/
@@ I Introduction
3) Comprehensive validation: The effectiveness of our approach is validated both in simulation and on real hardware, as well as across different robotic configurations.
Figure 2: Our method begins with a large-scale human motion dataset, which we retarget to the robot to extract an EE–elbow mapping dataset used to train our FiSTA network. After training, during human teleoperation we obtain the operator’s desired current EE pose via a VR headset. Given this target and a fixed-length history of past frames, FiSTA predicts the elbow position. We then augment the IK objective with an elbow-alignment cost and solve for the desired joint angles using Levenberg–Marquardt iterations. The resulting commands are sent to the robot’s low-level controller, yielding human-like motion.
@@ II-A Inverse Kinematics
The goal of IK is to compute joint configurations that realize a specified EE pose. Classical approaches pursue analytical (closed-form) solutions [2, 3, 4], which are highly efficient but hinge on restrictive assumptions about the manipulator’s kinematic structure. More recently, the field has shifted toward iterative optimization formulations [8, 9, 10, 19] that trade closed-form speed for modeling flexibility and can accommodate secondary objectives arising from kinematic redundancy, including geometric constraints [10, 20] and learned collision penalties [21]. Building on this line of work, our proposed HL-IK explicitly incorporates humanoid-likeness into the IK objective. While ensuring precise EE tracking, HL-IK further guide the arm configuration to resemble natural human arm postures, thereby achieving more anthropomorphic motion behavior.
@@ II-C Human-Like Motion Planning in IK
To encourage human-like arm configurations in IK, prior work largely takes two paths. Classical methods impose kinematic heuristics—e.g., swivel-angle parameterizations [26, 27] and physiology-inspired criteria [28]—but such hand-tuned rules generalize poorly across subjects and tasks. Recent data-driven method [29] learn wrist-pose→arm-angle mappings and then hard-wire them into closed-form IK; because this prior are entangled with analytic solvers, they transfer poorly to iterative numerical methods and restrict design flexibility. We instead learn an optimizer-agnostic elbow prior from large-scale human trajectories and encode it as a lightweight cost. Decoupled from any specific solver, this prior drops in as a generic objective term, integrating seamlessly with starndad numerical optimizers.
@@ III-A EE-Elbow Data Collection
We use AMASS [30] as the source of human motion trajectories, which provides consistent SMPL sequences with smooth kinematics and reliable pose quality. We retain six upper-limb joints—left/right shoulder, elbow, and EE—to represent the two arms. Following [17], we first estimate an upper-body shape parameter $\beta^{\prime}$ for the target robot via gradient-based fitting. Using $\beta^{\prime}$, we retarget each AMASS frame to the robot and build a frame-aligned dataset $\mathcal{D}_{\beta^{\prime}}$ paired with the original AMASS sequence $\mathcal{D}_{\beta}$.
@@ III-B FiLM-modulated Spatio-Temporal Attention Network for Elbow Prediction
Figure 4: Model Architecture. A GRU encodes the full history to produce temporal features, which are FiLM-conditioned using the next EE target. In parallel, an attention module computes spatial features from the last frame. The two streams are then concatenated to predict the elbow configuration.
@@ III-B3 Goal Modulator
The next-step EE target is mapped to FiLM-style feature-wise scale and shift, which affine-modulate the temporal summary and thus condition historical dynamics on the imminent goal.
@@ III-B4 Fusion and Output Head
The modulated temporal features are concatenated with the spatial summary and passed through an MLP regressor to produce the 7D elbow prediction.
@@ III-C2 Elbow Pose Cost
This elbow pose cost $\boldsymbol{c}_{\text{elbow}}$ enables the IK to track the given EE elbow to a certain extent.
@@ IV-A Model Evaluation
To balance efficiency and reliability, we adopt a two-stage training strategy: (1) All network comparison experiments, ablation studies, and hyperparameter searches are conducted on the ACCAD Dataset. (2) We then perform final training on the full dataset using the selected configuration. During comparison, we split the dataset by trajectory, with 90% for training and 10% for validation.
@@ IV-A1 Network Comparisons
We evaluated five representative architectures—MLP, LSTM, GRU, Transformer, and FiSTA—using the ACCAD subset, where each was trained with tuned hyperparameters under identical normalization and loss settings. The history length was fixed at 5 for all models. Validation errors are reported in Fig. 5. Among the candidates, FiSTA achieved the lowest validation error (MSE=0.001760), substantially outperforming all baselines. By contrast, the explicit architectural separation of spatial and temporal processing in FiSTA provides a stronger inductive bias than standard recurrent models or Transformers, which must learn these relationships from a less structured hidden state.
Figure 6: Ablation and sequence length comparison of FiSTA. Validation loss with different history lengths (1, 5, 8, 12, 16 frames) and ablation variants (without spatial, FiLM, or temporal).
@@ IV-A2 Tests on history length and ablation experiments on the FiSTA network
To rigorously evaluate our proposed model, we conducted a series of tests to analyze the impact of history sequence length and core architectural components on the final validation loss. First, we assessed the model’s performance with varying lengths of historical context, testing sequence lengths of 1, 5, 8, 12, and 16 frames. As shown in Fig. 6, the results clearly indicate an optimal contextual window. The model with a sequence length of 5 achieved the lowest validation loss, demonstrating maximum predictive precision. We observed that $L=5$ strikes a favorable balance: too short a context fails to provide sufficient dynamic information for inferring motion trends, while an excessively long context introduces irrelevant historical data, reducing the signal-to-noise ratio. To quantify the contribution of each component, we performed an ablation study on FiSTA (seq=5) (all deltas relative to the full model). Removing the Spatial Attention and Temporal modules increased the loss by 4.9% and 2.8%, respectively, indicating that both are important for building a robust spatiotemporal representation. The largest effect came from ablating the Goal Modulation module, which led to a 6.6% drop in accuracy. This module, implemented as a FiLM layer, conditions the GRU-encoded temporal features on the EE target. Its removal therefore highlights the critical role of FiLM-based goal conditioning in injecting target information into the temporal feature stream for this task.
@@
MLP 1.3104 $\pm$ 0.153 5.4896 $\pm$ 0.951 6.8000 $\pm$ 0.963 2.130
GRU 1.7565 $\pm$ 0.210 6.4981 $\pm$ 1.102 8.2547 $\pm$ 1.122 1.973
LSTM 1.5480 $\pm$ 0.185 4.3283 $\pm$ 0.824 5.8763 $\pm$ 0.842 2.197
Transformer 7.3642 $\pm$ 0.551 4.6888 $\pm$ 0.915 12.0530 $\pm$ 1.068 1.994
FiSTA (seq=1) 2.8091 $\pm$ 0.303 4.8272 $\pm$ 1.040 7.6363 $\pm$ 1.173 1.825
FiSTA (seq=5) 2.8665 $\pm$ 0.325 4.2175 $\pm$ 0.562 7.0840 $\pm$ 0.772 1.760
FiSTA_NoSpatial 2.0050 $\pm$ 0.126 3.9653 $\pm$ 0.679 5.9703 $\pm$ 0.698 1.846
FiSTA_NoTemporal 1.9993 $\pm$ 0.107 5.7616 $\pm$ 0.626 7.7609 $\pm$ 0.637 1.809
FiSTA_NoFiLM 2.1463 $\pm$ 0.072 9.0212 $\pm$ 3.291 11.1674 $\pm$ 3.303 1.876
FiSTA (seq=8) 3.0448 $\pm$ 0.339 4.2995 $\pm$ 0.745 7.3443 $\pm$ 0.954 1.923
FiSTA (seq=12) 3.1464 $\pm$ 0.380 5.4129 $\pm$ 1.188 8.5593 $\pm$ 1.405 1.902
FiSTA (seq=16) 3.1618 $\pm$ 0.368 4.2071 $\pm$ 0.658 7.3689 $\pm$ 0.878 1.926
Baseline IK (No NN) / 5.0818 $\pm$ 0.755 5.0818 $\pm$ 0.755 /
@@ IV-A3 Runtime Comparison
To assess the deployment potential of our model in practical applications, we conducted a comprehensive runtime tests on an NVIDIA RTX 4070 GPU. The total runtime for each model was calculated by averaging the runtime for one arm across the ACCAD trajectories. In this comparison, the total time is divided into two main stages as shown in Table I: (i) Preprocess, which includes the neural network’s input/output handling and the model inference itself. (ii) IK, which represents the time taken by the subsequent IK solver. Our optimal model, FiSTA (seq=5), achieved the best performance, though its overall running time ranked fourth among all the models. The final computation took 7.0840 (ms), which is well within the acceptable range.
@@ IV-A4 Conclusion
We set the default history length for FiSTA to $L=5$. The network was then trained on the entire dataset for a total of 1 hour and 50 minutes on an NVIDIA RTX 4090 GPU.
@@
Figure 8: The robot motions. Comparison of robot motions reconstructed by different methods: Ground Truth (green), HL-IK (red), and Baseline (blue). In each visualization, four spheres on each arm represent the shoulder, elbow, wrist, and $E$, respectively. The gray robot mesh is rendered from the ground-truth poses, while the colored skeletons illustrate the predicted trajectories of different methods.
@@ IV-B Comparative Analysis of Elbow Alignment
In $\mathbf{W}_{\mathrm{ee}}$ and $\mathbf{W}_{\mathrm{elbow}}$, the first $3{\times}3$ block weights translational residuals and the second weights rotational residuals; $\mathbf{I}_{d}$ denotes the $d{\times}d$ identity where $d$ is the joints number.
For quantitative evaluation, we use a new human-motion dataset (not for network training) - BMLhandball (649 trajectories, 183,806 steps) [40]. The trajectory generated by the original retargeting method serves as the reference, and we compare two IK solvers: (i) a baseline using only the EE pose as input (ii) our elbow-aligned method HL-IK. We conduct a comprehensive evaluation using four metrics: two that quantify arm-configuration similarity and two that EE tracking precision.
@@ IV-B4 EE Orientation Tracking Error
TABLE II: Comparison of Mean Metrics on the Full Dataset and Challenging Dateset
@@
Key Point Position Error (m) $5.944\times 10^{-2}\pm 0.00412$ $8.571\times 10^{-2}\pm 0.00942$ $5.977\times 10^{-2}\pm 0.00436$ $10.338\times 10^{-2}\pm 0.00812$
Line Angle Error (rad) $3.565\times 10^{-1}\pm 0.02438$ $5.515\times 10^{-1}\pm 0.08580$ $3.575\times 10^{-1}\pm 0.02190$ $6.791\times 10^{-1}\pm 0.06978$
EE Position Error (m) $4.575\times 10^{-3}\pm 0.00067$ $1.847\times 10^{-3}\pm 0.00028$ $5.405\times 10^{-3}\pm 0.00218$ $3.642\times 10^{-3}\pm 0.00483$
EE Orientation Error (rad) $8.110\times 10^{-4}\pm 0.00021$ $4.621\times 10^{-4}\pm 0.00008$ $10.728\times 10^{-4}\pm 0.00331$ $9.495\times 10^{-4}\pm 0.00141$
Based on the four metrics mentioned above, we conducted a comparative analysis on the BMLhandball’s [40] human retargeting dataset. First, we calculated the average error across all steps in the trajectory. The first two columns in Table. II reports the mean errors across all 183,806 steps. Each indicator in the table shows the average error for all steps. With elbow alignment, the arm’s key point position error and the line angle error are reduced of 30.6% and 35.4%, while the EE tracking accuracy drops slightly—an acceptable trade-off for teleoperation.
Since our aim is to probe configuration-tracking advantage in our HL-IK, we rank-ordered trajectories by the “Arm’s Key Point Position Error” achieved by the baseline method and selected the top 20% as “challenging trajectories” for further evaluation. We then compared the baseline with HL-IK on this subset. As reported in the last two columns in Table. II, and relative to the first two columns, the baseline shows large increases in both arm’s key point position error and line angel error, whereas HL-IK maintains nearly the same error levels as on the full set. On this subset, HL-IK reduces arm keypoint and direction errors by 42.2% and 47.4%, respectively. Fig. 8 visualizes this gap: the baseline (blue) deviates substantially, while HL-IK (red) remains closely aligned with the reference configuration.
Figure 9: Teleoperation on a physical robot. We deployed the HL-IK framework on the robot’s left arm (shown on the right side of each figure) and a standard IK solver without elbow-joint optimization on the right arm (shown on the left side of each figure). With HL-IK, the robot not only tracked the end effector but, without any direct sensing, also inferred and aligned with the human elbow position. As a result, its overall arm configuration closely matched the human’s, whereas the baseline IK achieved only end effector tracking and failed to reproduce configuration level similarity. Specifically, the higher the elbow joint height, the more apparent the difference.

### v3

@@ Abstract
Traditional IK methods for redundant humanoid manipulators emphasize end-effector (EE) tracking, frequently producing configurations that are valid mechanically but not human-like. We present Human-Like Inverse Kinematics (HL-IK), a lightweight IK framework that preserves EE tracking while shaping whole-arm configurations to appear human-like—without full-body sensing at runtime. The key idea is a learned elbow prior: using large-scale human motion data retargeted to the robot, we train a FiLM-modulated spatio-temporal attention network (FiSTA) to predict the next-step elbow pose from the EE target and a short history of EE–elbow states. This prediction is incorporated as a small residual alongside EE and smoothness terms in a standard Levenberg–Marquardt optimizer, making HL-IK a drop-in for numerical IK frameworks given a defined elbow frame and pose interface. Over 183k simulation steps, HL-IK reduces arm-similarity position and direction error by 30.6% and 35.4% on average, and by 42.2% and 47.4% on the most challenging trajectories. Hardware teleoperation on a robot distinct from simulation further confirms the gains in anthropomorphism. HL-IK is simple to integrate, adaptable across platforms via our pipeline, and adds minimal computation, enabling human-like motions for humanoid robots.
@@ I Introduction
3) Human-like IK (HL-IK) framework: We build an end-to-end HL-IK framework that (i) takes only EE targets as input, (ii) predicts a human-like elbow pose to resolve redundancy, and (iii) injects it as a residual cost into numerical IK, while keeping end-effector tracking as the primary objective and maintaining low runtime cost. Fig. 1 clearly illustrates the improved anthropomorphism achieved by our method.
Figure 2: Our method begins with a large-scale human motion dataset, which we retarget to the robot to extract an EE–elbow mapping dataset used to train our FiSTA network. After training, during human teleoperation we obtain the operator’s desired current EE pose via a VR. Given this target and a fixed-length history of past frames, FiSTA predicts the elbow position. We then augment the IK objective with an elbow-alignment cost and solve for the desired joint angles using Levenberg–Marquardt iterations. The resulting commands are sent to the robot’s low-level controller.
@@ II-A Inverse Kinematics
The goal of IK is to compute joint configurations that realize a specified EE pose. Classical approaches pursue analytical (closed-form) solutions [31, 38, 33], which are highly efficient but hinge on restrictive assumptions about the manipulator’s kinematic structure. More recently, the field has shifted toward iterative optimization formulations [27, 19, 30, 35] that trade closed-form speed for modeling flexibility and can accommodate secondary objectives arising from kinematic redundancy, including geometric constraints [30, 25] and learned collision penalties [26]. Building on this line of work, our proposed HL-IK explicitly incorporates humanoid-likeness into the IK objective. While ensuring precise EE tracking, HL-IK further guides the arm configuration to resemble natural human arm postures, thereby achieving more anthropomorphic motion behavior.
@@ II-C Human-Like Motion Planning in IK
To encourage human-like arm configurations in IK, prior work largely takes two paths. Classical methods impose kinematic heuristics—e.g., swivel-angle parameterizations [15, 17] and physiology-inspired criteria [39]—but such hand-tuned rules generalize poorly across subjects and tasks. We instead learn an optimizer-agnostic elbow prior from large-scale human trajectories and encode it as a lightweight cost. Decoupled from any specific solver, this prior drops in as a generic objective term, integrating seamlessly with standard numerical optimizers.
@@ III-A EE-Elbow Data Collection
We use AMASS [18] as the source of human motion trajectories, which provides consistent Skinned Multi-Person Linear model (SMPL) sequences with smooth kinematics and reliable pose quality. We retain six upper-limb joints—left/right shoulder, elbow, and EE—to represent the two arms. Following [12], we first estimate an upper-body shape parameter $\beta^{\prime}$ for the target robot via gradient-based fitting. Using $\beta^{\prime}$, we retarget each AMASS frame to the robot and build a frame-aligned dataset $\mathcal{D}_{\beta^{\prime}}$ paired with the original AMASS sequence $\mathcal{D}_{\beta}$.
@@ III-B FiLM-modulated Spatio-Temporal Attention Network for Elbow Prediction
Figure 4: Model Architecture. A GRU encodes the 5-frame history to produce temporal features, which are FiLM-conditioned using the next EE target. In parallel, an attention module computes spatial features from the last frame. The two streams are then concatenated to predict the elbow configuration.
@@ III-B3 Goal Modulator
The next-step EE target is mapped to Feature-wise Linear Modulation (FiLM)-style [24] feature-wise scale and shift, which affine-modulate the temporal summary and thus condition historical dynamics on the imminent goal.
@@ III-B4 Fusion and Output Head
The modulated temporal features are concatenated with the spatial summary and passed through an Multi-Layer Perceptron(MLP) regressor to produce the 7D elbow prediction.
@@ III-C2 Elbow Pose Cost
This elbow pose cost $\boldsymbol{c}_{\text{elbow}}$ enables the IK to track the given predicted elbow pose to a certain extent.
@@ IV-A Model Evaluation
To balance efficiency and reliability, we adopt a two-stage training strategy: (1) All network comparison experiments, ablation studies, and hyperparameter searches are conducted on the ACCAD Dataset. (2) We then perform final training on the full dataset using the selected configuration. During comparison, we split the dataset by trajectory, with 90% for training and 10% for validation. All experiments are repeated with four random seeds (42, 123, 789, 1024) to ensure statistical reliability, and we report mean $\pm$ standard deviation.
@@ IV-A1 Network Comparisons
We evaluated five representative architectures—MLP, Long Short-Term Memory (LSTM), Gated Recurrent Unit (GRU), Transformer, and FiSTA—using the ACCAD subset, where each was trained with tuned hyperparameters under identical normalization and loss settings. The network is trained by minimizing the Mean Squared Error (MSE) and all baseline models share the same input-output specification as FiSTA. The history length was fixed at 5 for all models. Validation errors are reported in Fig. 5. Among the candidates, FiSTA achieved the lowest validation error (MSE=0.001781 $\pm$ 0.000014), substantially outperforming all baselines. By contrast, the explicit architectural separation of spatial and temporal processing in FiSTA provides a stronger inductive bias than standard recurrent models or Transformers, which must learn these relationships from a less structured hidden state.
Figure 6: Ablation and sequence length comparison of FiSTA. Validation loss with different history lengths (1, 5, 8, 12, 16 frames) and ablation variants (without spatial, FiLM, or temporal). Shaded regions indicate standard deviation across four random seeds.
@@ IV-A2 Tests on history length and ablation experiments on the FiSTA network
To rigorously evaluate our proposed model, we conducted a series of tests to analyze the impact of history sequence length and core architectural components on the final validation loss. First, we assessed the model’s performance with varying lengths of historical context, testing sequence lengths of 1, 5, 8, 12, and 16 frames. As shown in Fig. 6, the results clearly indicate an optimal contextual window. The model with a sequence length of 5 achieved the lowest validation loss, demonstrating maximum predictive precision. We observed that $L=5$ strikes a favorable balance: too short a context fails to provide sufficient dynamic information for inferring motion trends, while an excessively long context introduces irrelevant historical data, reducing the signal-to-noise ratio. To quantify the contribution of each component, we performed an ablation study on FiSTA (seq=5) (all deltas relative to the full model). Removing the Spatial Attention and Temporal modules increased the loss by 4.5% and 2.3%, respectively, indicating that both are important for building a robust spatiotemporal representation. The largest effect came from ablating the Goal Modulation module, which led to a 6.0% drop in accuracy. This module, implemented as a FiLM layer, conditions the GRU-encoded temporal features on the EE target. Its removal therefore highlights the critical role of FiLM-based goal conditioning in injecting target information into the temporal feature stream for this task.
@@
MLP 1.3104 $\pm$ 0.153 5.4896 $\pm$ 0.951 6.8000 $\pm$ 0.963 2.106 $\pm$ 0.029
GRU 1.7565 $\pm$ 0.210 6.4981 $\pm$ 1.102 8.2547 $\pm$ 1.122 2.025 $\pm$ 0.045
LSTM 1.5480 $\pm$ 0.185 4.3283 $\pm$ 0.824 5.8763 $\pm$ 0.842 2.040 $\pm$ 0.055
Transformer 7.3642 $\pm$ 0.551 4.6888 $\pm$ 0.915 12.0530 $\pm$ 1.068 2.390 $\pm$ 0.135
FiSTA (seq=1) 2.8091 $\pm$ 0.303 4.8272 $\pm$ 1.040 7.6363 $\pm$ 1.173 1.890 $\pm$ 0.030
FiSTA (seq=5) 2.8665 $\pm$ 0.325 4.2175 $\pm$ 0.562 7.0840 $\pm$ 0.772 1.781 $\pm$ 0.014
FiSTA_NoSpatial 2.0050 $\pm$ 0.126 3.9653 $\pm$ 0.679 5.9703 $\pm$ 0.698 1.862 $\pm$ 0.037
FiSTA_NoTemporal 1.9993 $\pm$ 0.107 5.7616 $\pm$ 0.626 7.7609 $\pm$ 0.637 1.822 $\pm$ 0.020
FiSTA_NoFiLM 2.1463 $\pm$ 0.072 9.0212 $\pm$ 3.291 11.1674 $\pm$ 3.303 1.889 $\pm$ 0.008
FiSTA (seq=8) 3.0448 $\pm$ 0.339 4.2995 $\pm$ 0.745 7.3443 $\pm$ 0.954 1.890 $\pm$ 0.038
FiSTA (seq=12) 3.1464 $\pm$ 0.380 5.4129 $\pm$ 1.188 8.5593 $\pm$ 1.405 1.935 $\pm$ 0.010
FiSTA (seq=16) 3.1618 $\pm$ 0.368 4.2071 $\pm$ 0.658 7.3689 $\pm$ 0.878 1.923 $\pm$ 0.018
Jacobian-based numerical IK / 5.0818 $\pm$ 0.755 5.0818 $\pm$ 0.755 /
@@ IV-A3 Runtime Comparison
To assess the deployment potential of our model in practical applications, we conducted a comprehensive runtime test on an NVIDIA RTX 4070 GPU. The total runtime for each model was calculated by averaging the runtime for one arm across the ACCAD trajectories. In this comparison, the total time is divided into two main stages as shown in Table I: (i) Preprocess, which includes the neural network’s input/output handling and the model inference itself. (ii) IK, which represents the time taken by the subsequent IK solver. Our optimal model, FiSTA (seq=5), achieved the best performance, though its overall running time ranked fourth among all the models. The final computation takes 7.0840 ms per step, compared to 5.0818 ms for the Jacobian-based numerical IK, introducing only an additional  2.0022 ms overhead. Even with this overhead, the runtime still supports a control frequency of approximately 141.2 Hz, which satisfies common real-time control requirements.
@@ IV-A4 Conclusion
We set the default history length for FiSTA to $L=5$. The network was then trained on the entire dataset for a total of 1 hour and 50 minutes on an NVIDIA RTX 4090 GPU. This final trained model is used for all subsequent elbow-alignment experiments in Section IV-B and IV-C.
@@
Figure 8: The robot motions. Comparison of robot motions reconstructed by different methods: Ground Truth (green), HL-IK (red), and swivel-angle–based Baseline (blue). In each visualization, four spheres on each arm represent the shoulder, elbow, wrist, and $E$, respectively. The gray robot mesh is rendered from the ground-truth poses, while the colored skeletons illustrate the predicted trajectories of different methods.
@@ IV-B Comparative Analysis of Elbow Alignment
In $\mathbf{W}_{\mathrm{ee}}$ and $\mathbf{W}_{\mathrm{elbow}}$, the first $3{\times}3$ block weights translational residuals and the second weights rotational residuals; $\mathbf{I}_{d}$ denotes the $d{\times}d$ identity where $d$ is the number of joints.
For quantitative evaluation, we use a new human-motion dataset (not for network training) - BMLhandball (649 trajectories, 183,806 steps) [13]. The trajectory generated by the original retargeting method serves as the reference. We compare two IK solvers: (i) a swivel-angle–based baseline, following the method in [17], and (ii) our elbow-aligned approach, HL-IK. We conduct a comprehensive evaluation using four metrics: two that quantify arm-configuration similarity and two that EE tracking precision.
@@ IV-B4 EE Orientation Tracking Error
TABLE II: Comparison of Mean Metrics on the Full Dataset and Challenging Dataset
@@
Key Point Position Error (m) $\mathbf{(5.944\pm 0.412)\times 10^{-2}}$ $(8.571\pm 0.942)\times 10^{-2}$ $\mathbf{(5.977\pm 0.436)\times 10^{-2}}$ $(10.338\pm 0.812)\times 10^{-2}$
Line Angle Error (rad) $\mathbf{(3.565\pm 0.244)\times 10^{-1}}$ $(5.515\pm 0.858)\times 10^{-1}$ $\mathbf{(3.575\pm 0.219)\times 10^{-1}}$ $(6.791\pm 0.698)\times 10^{-1}$
EE Position Error (m) $(4.575\pm 0.670)\times 10^{-3}$ $\mathbf{(1.847\pm 0.280)\times 10^{-3}}$ $(5.405\pm 2.180)\times 10^{-3}$ $\mathbf{(3.642\pm 4.830)\times 10^{-3}}$
EE Orientation Error (rad) $(8.110\pm 2.100)\times 10^{-4}$ $\mathbf{(4.621\pm 0.800)\times 10^{-4}}$ $(10.728\pm 33.10)\times 10^{-4}$ $\mathbf{(9.495\pm 14.10)\times 10^{-4}}$
Based on the four metrics mentioned above, we conducted a comparative analysis on the BMLhandball’s [13] human retargeting dataset. First, we calculated the average error across all steps in the trajectory. The first two columns in Table II report the mean errors across all 183,806 steps. Each indicator in the table shows the average error for all steps.The arm keypoint position error and line-angle error are reduced by 30.6% and 35.4% compared with the baseline. Meanwhile, EE tracking accuracy decreases only slightly, with an increase of 2.728 mm in EE position error and 0.003489 rad in EE orientation error—an acceptable trade-off for teleoperation.
Since our aim is to probe configuration-tracking advantage in our HL-IK, we rank-ordered trajectories by the “Arm’s Key Point Position Error” achieved by the baseline method and selected the top 20% as “challenging trajectories” for further evaluation. We then compared the baseline with HL-IK on this subset. As reported in the last two columns in Table  II, and relative to the first two columns, the baseline shows large increases in both arm’s key point position error and line angle error, whereas HL-IK maintains nearly the same error levels as on the full set. On this subset, HL-IK reduces arm keypoint and direction errors by 42.2% and 47.4%, respectively. Fig. 8 visualizes this gap: the baseline (blue) deviates substantially, while HL-IK (red) remains closely aligned with the reference configuration.
Figure 9: Teleoperation on a physical robot. We deployed the HL-IK framework on the robot’s left arm (shown on the right side of each figure) and a Jacobian-based numerical IK to the right arm, IK solver without elbow-joint optimization on the right arm (shown on the left side of each figure). With HL-IK, the robot not only tracked the end effector but, without any direct sensing, also inferred and aligned with the human elbow position. As a result, its overall arm configuration closely matched the human’s, whereas the standard IK achieved only end effector tracking and failed to reproduce configuration level similarity. Specifically, the higher the elbow joint height, the more apparent the difference.